# Blockchain with proof of quantum work

### v2

@@ Abstract
We propose a blockchain architecture in which mining requires a quantum computer. The consensus mechanism is based on proof of quantum work, a quantum-enhanced alternative to traditional proof of work that leverages quantum supremacy to make mining intractable for classical computers. We have refined the blockchain framework to incorporate the probabilistic nature of quantum mechanics, ensuring stability against sampling errors and hardware inaccuracies. To validate our approach, we implemented a prototype blockchain on four D-WaveTM quantum annealing processors geographically distributed within North America, demonstrating stable operation across hundreds of thousands of quantum hashing operations. Our experimental protocol follows the same approach used in the recent demonstration of quantum supremacy [1], ensuring that classical computers cannot efficiently perform the same computation task. By replacing classical machines with quantum systems for mining, it is possible to significantly reduce the energy consumption and environmental impact traditionally associated with blockchain mining while providing a quantum-safe layer of security. Beyond serving as a proof of concept for a meaningful application of quantum computing, this work highlights the potential for other near-term quantum computing applications using existing technology.
@@ I Introduction
Traditional blockchains use consensus algorithms like proof of work (PoW) and proof of stake (PoS) to validate transactions. PoW relies upon the significant computational effort to generate valid blocks. Distributed computational power ensures that no user can control the blockchain evolution. Introducing invalid transactions, such as double-spending, requires of the order of 50% of the community compute power, which is financially prohibitive due to the substantial equipment and electricity costs. While PoW effectively enhances network security, it also leads to significant energy consumption. Bitcoin mining alone consumes an estimated 175.87 terawatt-hours (TWh) in 2024, surpassing the yearly electricity usage of entire countries, including Sweden [12, 13]. This high energy consumption raises environmental concerns.
@@
As blockchain technology advances, researchers and technologists are investigating innovative approaches to overcome its limitations and ensure its long-term viability. One exciting avenue is the integration of quantum computing with blockchain systems, a combination that could transform the field by offering unmatched security, scalability, and computational power. This article introduces a blockchain architecture that incorporates a proof of quantum work (PoQ) consensus mechanism, where quantum computers are required to validate transactions. The central concept is to harness quantum supremacy to mine problems intractable to classical computers. Quantum supremacy is a term introduced by John Preskill [16] that refers to a scenario in which a quantum computer performs a task that classical computers cannot accomplish with reasonable resources. While there could be various ways to leverage quantum supremacy within a blockchain, our proposal builds on Bitcoin’s architecture with minimal modifications, taking advantage of its proven reliability and robustness.
Unlike some previous quantum blockchain proposals [17, 18, 19, 20, 21, 22, 23, 24, 25, 26], which often assume access to fault-tolerant quantum computing or quantum teleportation, our approach integrates both classical and quantum computation to leverage quantum supremacy with existing and near-term QPUs. Our proposal is also different from blockchains and hashes that have been proposed to exploit the potential of near-term boson-sampling [27, 28, 29]. We introduce a hierarchy of practical witness methods, new consensus mechanisms robust to noise, and implement a proof-of-concept experiment using a distributed network of quantum computers.
Although claims of supremacy have been made in Boson sampling, these have been challenged [30, 31]. Currently, Ref. [1] is the only demonstration of quantum supremacy not based on random-circuit sampling [32, 33], and thus the only demonstration of quantum supremacy relevant to this work (see Appendix A). The unitary dynamics used for hashing remains intractable despite recent developments in algorithmic methods [34, 35, 36].
Using a small-scale prototype, we demonstrate this blockchain’s functionality on four generally-accessible D-Wave AdvantageTM and Advantage2TM quantum processing units (QPUs) at different geographic locations across North America. By carefully accounting for the probabilistic nature of quantum computation and potential errors within Bitcoin’s architecture, we achieve stable operation across thousands of mining processes distributed among hundreds of stakeholders. This marks the first successful demonstration of an important real-world application that directly leverages quantum supremacy, operating seamlessly across an internationally distributed network of quantum computers.
@@ II Classical Blockchain Architecture
Transactions are recorded in blocks linked sequentially in a chain, forming a continuous and immutable ledger. A Bitcoin block consists of a block header and a list of transactions (see Figure 1 for illustration). The block header acts as a metadata section and contains essential elements such as a timestamp, a Merkle root (a hierarchical summary of all transactions in the block), the nonce, and the hash of the previous block. A hash is a unique, fixed-length binary string generated by processing data through a cryptographic algorithm. It acts as a digital fingerprint, ensuring data integrity by making even the slightest change to the input data result in a completely different hash. The nonce is an arbitrary 32-bit number adjusted by the miners to solve the cryptographic puzzle required for block validation. Miners achieve this by repeatedly combining the transaction data with the nonce and passing it through a classical hashing algorithm, which generates an ${\cal N_{H}}$-bit hash ${\cal H}$ (${\cal N_{H}}=256$ for Bitcoin). Adding a block to the blockchain begins with miners competing to find a nonce that produces a hash value below a specific target ${\cal H}<{\cal T}_{n}$, with $n$ indexing the block. The inequality requires that the first ${\cal N}_{\rm zeros}$ bits of the hash be zero, where
@@ III Classical Hashing
IV Quantum Hashing
Figure 2: Illustration of the quantum hash generation and its use as proof of work for block security.
@@ IV Quantum Hashing
The last requirement is essential when the hash is used in a blockchain with a PoQ-based consensus mechanism. It ensures that only quantum computers can perform mining, thereby preventing energy-intensive classical mining. The probabilistic nature of quantum hash values must be carefully considered in the design and application of hash functions, particularly in systems like blockchain, where reliable behavior is essential. This issue, along with potential solutions, will be discussed in greater detail in subsequent sections. Since existing standard classical hashes already fulfill the first five classical requirements, by carefully combining a classical hash function with a quantum operation, the focus can shift to achieving the last requirement.
We start by introducing the general concept without delving into specific implementation details. Let $\mathcal{M}$ represent a message of arbitrary length. (For clarity and consistency, we use calligraphic symbols for blockchain-related quantities, while standard symbols are reserved for all other variables, including those characterizing the quantum evolution.) A hash function is defined as a unique transformation of $\mathcal{M}$ into a hash value ${\cal H}(\mathcal{M})$, with a fixed length of $\mathcal{N}_{\mathcal{H}}$ bits. This transformation is performed in four steps (see Fig. 2 for visualization):
@@
The 4-stage process outlined includes both classical and quantum computation. Although the classical part likely includes non-trivial operations such as cryptographic hashing, we assume that the quantum computation is significantly more costly (and time-consuming) than the sum of classical operations, which is true in our experimental implementation.
@@
The postprocessing step then aims to convert $P_{\mathcal{M}}$ into a unique bit string of predetermined length $\mathcal{N}_{\mathcal{H}}$, ensuring robustness to uncertainty in the experiment whilst capturing non-trivial statistical properties. In other words, the resulting quantum hash is a fixed-length bit string that uniquely identifies the probability distribution $P_{\mathcal{M}}$. A complete characterization of $P_{\mathcal{M}}$ for a system with $N_{Q}$ qubits requires $12^{N_{Q}}{-}$ 1 real parameters. This characterization can be achieved by determining the probability of each state in the computational basis or by calculating all expectation values of single- and multi-qubit operators up to order $N_{Q}$, i.e., $\langle\sigma^{z}_{i}\rangle$, $\langle\sigma^{z}_{i}\sigma^{z}_{j}\rangle$, $\langle\sigma^{z}_{i}\sigma^{z}_{j}\sigma^{z}_{k}\rangle$, etc. In practice, for quantum systems with local interactions, it is often sufficient to consider a limited number of dominant k-local expectations within the correlation length to characterize the distribution with acceptable accuracy. This simplification significantly reduces the complexity of the problem, particularly when the primary goal is to ensure collision resistance in the resulting quantum hash.
@@
Each element of the feature vector $V$ is a real number $-1{\leq}V_{\alpha}{\leq}1$. To generate an ${\cal N_{H}}$-bit hash we create ${\cal N_{H}}$ witnesses as linear combinations of $V_{\beta}$
@@
where $G$ is an ${\cal N_{H}}{\times}N_{V}$ matrix with $O(1)$ elements such that $-1{\leq}W_{\alpha}{\leq}1$. We turn each element of the witness vector $W$ into a binary number using (6). Each equation $W_{\alpha}=W_{0\alpha}$ defines a hyperplane in the vector space made of $V$’s and each hash bit $\mathcal{H}_{\alpha}$ indicates on which side of the hyperplane the observed $V$ lies. The threshold value $W_{0\alpha}$ can be taken to be zero or as some simple function of the experimental parameters. Robustness of the hash bits towards errors and spoofing depends on the choice of the hyperplanes (see Appendix C.5). The signs on hyperplane projections can be chosen uniformly at random as a function of the message (as in our proof of concept) to guarantee a uniform distribution of bit strings that is not a reversible function of other experimental parameters, providing pre-image protection.
When qubit readout is performed in a single basis, it is, in principle, possible to identify a classical system that can replicate the same distribution without entanglement. While finding such a classical system may be practically infeasible due to quantum supremacy, its theoretical existence cannot be ruled out. Indeed, any proof of entanglement, whether through an entanglement measure or a witness, requires access to information beyond single-basis measurements of the quantum states. This additional information could be obtained by performing measurements in multiple bases or by examining the system’s response to perturbations (e.g., susceptibility measurements). In the absence of more complex measurement, one may resort to susceptibility measurements to protect further against classical spoofing.
@@ V.1 Probabilistic Validation
We now develop a model of probabilistic validation assuming independent and Gaussian distributed witness statistics, which is consistent with witnesses being determined by a sum of many weakly correlated statistics, and proves to be a good approximation in experiments as shown later. We introduce a confidence measure that quantifies the accuracy of hash generation and validation relative to an ideal, error-free scenario. The proposed confidence measure can serve as a principled heuristic for practical applications, independent of the underlying assumptions
@@
$P(W_{\alpha})={1\over\sqrt{2\pi}\delta W_{\alpha}}e^{-(W_{\alpha}-\bar{W}_{\alpha})^{2}/2\delta W_{\alpha}^{2}}.$ (11)
@@
[…]
@@ V.2 Redefining Chainwork
If the work remains constant, then one can assign ±$\pm 1$ to valid/invalid blocks, we call this basic +/- Chainwork. In the case of a tie, the first path evaluated is preferred, following the same rule as in Bitcoin. This adjustment effectively resolves the issue by turning a hard validation fork to a soft validation fork that is resolved subject to a delay. If a user falls behind by two blocks, the majority compute power of miners will produce a third block, causing the strongest chain to advance by three blocks in agreement with the majority.
@@
VI Experimental Implementation
Experimental realization of a quantum blockchain requires access to a quantum computer capable of solving complex problems where quantum supremacy has been demonstrated or is potentially achievable. The generated distribution must exhibit sufficient structure to allow observables with nontrivial expectation values that can be measured reproducibly across multiple QPUs. Therefore, techniques such as boson sampling or random circuit sampling which cannot be formulated as efficient decision problems, are not suitable for this application (see Appendix A). Coarse-grained boson sampling has been proposed as a means to implement blockchains [28, 29], but to our knowledge there is no experimental implementation (see Appendix A.1). Currently, Ref. [1] is the only experimental demonstration of quantum supremacy that meets the requirement of efficient cross-validation. Therefore, we base our experimental demonstration on the same protocol. Since Ref. [1] has already established agreement with quantum mechanics and the infeasibility of classical simulation, our primary focus here is to demonstrate that multiple QPUs can generate and validate hashes, ensuring the blockchain operates reliably with high efficiency and small transaction delays.
Figure 3: Operation of an example blockchain with $50$ miners using four publicly accessible Advantage and Advantage2 QPUs. The mining difficulty was set to $\mathcal{N}_{\rm zeros}=32$ with basic +/-1 Chainwork defined in (18). Orange blocks show soft (resolved) forks, these blocks are included in no strongest chains (have been rejected by all miners). Blue blocks are immutable, included in all strongest chains (agreed upon by all miners). Gray and black blocks are in contention: common to some (but not all) strongest chains. Only black blocks are candidates for maximum Chainwork (are being mined) and have the potential for further branching. Blocks are labeled by the order of broadcast with the genesis block on the left and most recently mined block on the right. Either of the paths terminating in black nodes might be extended, but the orange path is no longer a candidate for the strongest chain. Blocks 10, 11, 12 and 13 contain delayed transactions, that require further mining (fork resolution) in order to be confirmed. A majority of miners believe node 13 has maximum Chainwork, and in this experiment all miners consolidate on this branch as the chain develops (see Figure 4).
@@ VI Experimental Implementation
Figure 5: Mining efficiency, i.e., the fraction of blocks joining the strongest chain. The Chainwork uses either the simple +/-1 weighting, or the confidence-based weighting with $\mathcal{N}_{\rm max}=1,$ $2$ or $3$ with a constant $\delta W_{\alpha}$ extracted from experimental data. Confidence-weighted strongest chains are more efficient. The mining process is accelerated without impact on statistical outcomes as described in Appendix  D.3. Results are obtained with resampling from experimentally-parameterized witness distributions to accelerate analysis, each point an average on 16 chains of length 512-1024. Resampling statistics are in agreement with full-experiment efficiencies reported in Figures 4 and 10. Further experimental details and statistics are contained in Appendix D.
The experiments are conducted on four D-Wave Advantage and Advantage2 QPUs. Each QPU comprises a network of coupled qubits that undergo continuous evolution governed by a time-dependent Hamiltonian
@@
The quantum problem we solve for hash generation is the coherent quenching of spin glasses through quantum phase transitions. We embed cubic lattices of size $4{\times}4{\times}$ 4 × 4 dimers (128 qubits) on each of four different internationally distributed QPUs, supplementary results for $18{\times}18$ dimerized Biclique models (72 qubits) are limited to Appendix D. The critical dynamics ensure that by adjusting a single parameter (annealing time) different QPUs can produce consistent results despite variations in their energy scales [1]. All qubit biases are set to zero, $h_{i}=0$, and couplers are selected as a function of the input message $J_{ij}=\Theta(\mathcal{M})$. This is done in two steps: First the message goes through a classical hash, e.g., SHA-256 in our case. Then the classical hash is used as a seed in a pseudo-random number generator that generates the experimental parameters ($J$ only, in our proof of concept). The classical hash is also used as the cryptographic identifier of the block in the block header of the next block. The quantum hash is only used for PoQ. This way, we can choose ${\cal N_{H}}={\cal N}_{\rm zeros}$ because nonzero bits in the quantum hash have no significance. Using a spin glass Hamiltonian for quantum evolution ensures that classical collision resistance carries over to the quantum hash. It is extremely unlikely that two spin glass Hamiltonians with entirely different randomly chosen parameters collide in the resulting high dimensional distribution. Furthermore, the projection matrix $G(\mathcal{M})$ is also randomly generated based on the message $\mathcal{M}$, making collisions even more improbable.
@@
Fig. 3 presents the initial-stage operation of a blockchain using four publicly accessible D-Wave Advantage and Advantage2 QPUs, involving 50 miners with $\mathcal{N}_{\rm zeros}=32$. The visualization captures the dynamic evolution of the blockchain, highlighting block statuses and mining activity. Blocks on the (majority) strongest chain are arranged chronologically from left to right, with the genesis block on the far left and the most recently mined block on the far right. Different block colors indicate their status within the consensus process: Blue blocks (immutable) are included in all strongest chains and are universally agreed upon. Orange blocks (soft forks) have been rejected by all miners and are not included in any of the strongest chains. Gray and black blocks appear in some but not all strongest chains, reflecting temporary disagreement among miners. Only black blocks are actively being mined, meaning they have the potential for further branching. This illustrates how consensus emerges dynamically as blocks are proposed and either accepted or rejected by the network.
@@
In Fig. 4, the blockchain is shown after 219 block broadcasts, with the genesis block repositioned centrally for clarity. By this stage, approximately 70% of proposed blocks have become immutable (blue), aligning well with theoretical expectations based on the bootstrapping methodology (see Appendix D.3). The remaining blocks continue to be evaluated, with only those on the strongest chains persisting in the blockchain ledger. This simulation demonstrates the robustness of the blockchain framework, showcasing its ability to maintain stability and achieve miner consensus despite the probabilistic nature of quantum computation. We are able to use statistical properties of the Chainwork described in Section D.2 to eliminate failed mining experiments, and create the blockchain with only $50\times 219$ total QPU programmings. Blockchains with more miners, and over $1000$ block broadcasts, are shown in Appendices.

### v3

@@ Abstract
We propose a blockchain architecture in which mining requires a quantum computer. The consensus mechanism is based on proof of quantum work, a quantum-enhanced alternative to traditional proof of work that leverages beyond-classical quantum computation to make mining intractable for classical computers. We have refined the blockchain framework to incorporate the probabilistic nature of quantum mechanics, ensuring stability against sampling errors and hardware inaccuracies. To validate our approach, we implemented a prototype blockchain on four D-WaveTM quantum annealing processors geographically distributed within North America, demonstrating stable operation across hundreds of thousands of quantum hashing operations. Our sampling protocol follows the same approach used in the recent demonstration of beyond-classical computation [1], ensuring that classical computers cannot efficiently perform the same computation task. By replacing classical machines with quantum systems for mining, it is possible to significantly reduce the energy consumption and environmental impact traditionally associated with blockchain mining while providing a quantum-safe layer of security. Beyond serving as a proof of concept for a meaningful application of quantum computing, this work highlights the potential for other near-term quantum computing applications using existing technology.
@@ I Introduction
Traditional blockchains use consensus algorithms like proof of work (PoW) and proof of stake (PoS) to validate transactions. PoW relies upon the significant computational effort to generate valid blocks. Distributed computational power ensures that no user can control the blockchain evolution. Introducing invalid transactions, such as double-spending, requires of the order of 50% of the community compute power, which is financially prohibitive due to the substantial equipment and electricity costs. While PoW effectively enhances network security, it also leads to significant energy consumption. Bitcoin mining alone consumed an estimated 175.87 terawatt-hours (TWh) in 2024, surpassing the yearly electricity usage of entire countries, including Sweden [12, 13]. This high energy consumption raises environmental concerns.
@@
As blockchain technology advances, researchers and technologists are investigating innovative approaches to overcome its limitations and ensure its long-term viability. One exciting avenue is the integration of quantum computing with blockchain systems, a combination that could transform the field by offering unmatched security, scalability, and computational power. This article introduces a blockchain architecture that incorporates a proof of quantum work (PoQ) consensus mechanism, where quantum computers are required to validate transactions. The central concept is to harness beyond-classical quantum computation to mine problems intractable to classical computers. Quantum supremacy is a term introduced by John Preskill [16] that refers to a scenario in which a quantum computer performs a task that classical computers cannot accomplish with reasonable resources. While there could be various ways to leverage this quantum advantage within a blockchain, our proposal builds on Bitcoin’s architecture with minimal modifications, taking advantage of its proven reliability and robustness.
Unlike some previous quantum blockchain proposals [17, 18, 19, 20, 21, 22, 23, 24, 25, 26], which often assume access to fault-tolerant quantum computing or quantum teleportation, our approach integrates both classical and quantum computation to leverage quantum advantage with existing and near-term QPUs. Our proposal is also different from blockchains and hashes that have been proposed to exploit the potential of near-term boson-sampling [27, 28, 29]. We introduce a hierarchy of practical witness methods, new consensus mechanisms robust to noise, and implement a proof-of-concept demonstration using a distributed network of quantum computers.
Although claims of beyond-classical computation have been made in Boson sampling, these have been challenged [30, 31]. Currently, Ref. [1] is the only demonstration of beyond-classical computation not based on random-circuit sampling [32, 33], and thus the only demonstration of quantum advantage relevant to this work (see Appendix A). The unitary dynamics used for hashing remains intractable despite recent developments in algorithmic methods [34, 35, 36].
Using a small-scale prototype, we demonstrate this blockchain’s functionality on four generally-available D-Wave AdvantageTM and Advantage2TM quantum processing units (QPUs) at different geographic locations across North America. By carefully accounting for the probabilistic nature of quantum computation and potential errors within Bitcoin’s architecture, we achieve stable operation across thousands of mining processes distributed among hundreds of stakeholders. This marks the first successful demonstration of an important real-world application that directly leverages beyond-classical computation, operating seamlessly across an internationally distributed network of quantum computers.
@@ II Classical Blockchain Architecture
[…]
@@ III Classical Hashing
IV Hybrid Classical-Quantum Hashing
Figure 2: Illustration of the hybrid classical-quantum hash generation and its use as proof of work for block security.
@@ IV Hybrid Classical-Quantum Hashing
The last requirement is essential when the hash is used in a blockchain with a PoQ-based consensus mechanism. It ensures that only quantum computers can perform mining, thereby preventing energy-intensive classical mining. The probabilistic nature of quantum hash values must be carefully considered in the design and application of hash functions, particularly in systems like blockchain, where reliable behavior is essential. This issue, along with potential solutions, will be discussed in greater detail in subsequent sections. Rather than requiring our quantum evolution to provably guarantee cryptographic properties, it is convenient to exploit a hybrid approach with one classical cryptographic hash used to define the quantum-computation parameters. Since existing classical cryptographic hashes already fulfill the first five classical requirements the focus can shift to achieving the last requirement.
[…]
@@
The 4-stage process outlined includes both classical and quantum computation. Although the classical part likely includes non-trivial operations such as cryptographic hashing, it is straightforward to demonstrate the quantum computation costs (time, energy or dollars) are significantly larger than the combination of classical operations, as we later quantify in context of our demonstration.
@@
The postprocessing step then aims to convert $P_{\mathcal{M}}$ into a unique bit string of predetermined length $\mathcal{N}_{\mathcal{H}}$, ensuring robustness to uncertainty in the experiment whilst capturing non-trivial statistical properties. In other words, the resulting quantum hash is a fixed-length bit string that uniquely identifies the probability distribution $P_{\mathcal{M}}$. A complete characterization of $P_{\mathcal{M}}$ for a system with $N_{Q}$ qubits requires $2^{N_{Q}}{-}1$ real parameters. This characterization can be achieved by determining the probability of each state in the computational basis or by calculating all expectation values of single- and multi-qubit operators up to order $N_{Q}$, i.e., $\langle\sigma^{z}_{i}\rangle$, $\langle\sigma^{z}_{i}\sigma^{z}_{j}\rangle$, $\langle\sigma^{z}_{i}\sigma^{z}_{j}\sigma^{z}_{k}\rangle$, etc. In practice, for quantum systems with local interactions, it is often sufficient to consider a limited number of dominant k-local expectations within the correlation length to characterize the distribution with acceptable accuracy. This simplification significantly reduces the complexity of the problem, particularly when the primary goal is to ensure collision resistance in the resulting quantum hash.
@@
Each element of the feature vector $V$ is a real number $1{\leq}V_{\alpha}{\leq}1$. To generate an $\cal N_{H}$}-bit hash we create $\cal N_{H}$} witnesses as linear combinations of $V_{\beta}$
@@
where $G$ is an ${\cal N_{H}}{\times}N_{V}$ matrix with $O(1)$ elements such that $1{\leq}W_{\alpha}{\leq}1$. We turn each element of the witness vector $W$ into a binary number using (6). Each equation $W_{\alpha}=W_{0\alpha}$ defines a hyperplane in the vector space made of $V$’s and each hash bit $\mathcal{H}_{\alpha}$ indicates on which side of the hyperplane the observed $V$ lies. The threshold value $W_{0\alpha}$ can be taken to be zero or as some simple function of the experimental parameters. Robustness of the hash bits towards errors and spoofing depends on the choice of the hyperplanes (see Appendix C.5). The signs on hyperplane projections can be chosen uniformly at random as a cryptographic function of the message (as in our proof of concept) to guarantee a uniform distribution of bit strings that is not a reversible function of other experimental parameters, providing pre-image protection.
When qubit readout is performed in a single basis, it is, in principle, possible to identify a classical system that can replicate the same distribution without entanglement. While finding such a classical system may be practically infeasible due to beyond-classical computation, its theoretical existence cannot be ruled out. Indeed, any proof of entanglement, whether through an entanglement measure or a witness, requires access to information beyond single-basis measurements of the quantum states. This additional information could be obtained by performing measurements in multiple bases or by examining the system’s response to perturbations (e.g., susceptibility measurements). In the absence of more complex measurement, one may resort to susceptibility measurements to protect further against classical spoofing.
@@ V.1 Probabilistic Validation
We now develop a model of probabilistic validation assuming independent and Gaussian distributed witness statistics, which is consistent with witnesses being determined by a sum of many weakly correlated statistics, and proves to be a good approximation in demonstrations as shown later. We introduce a confidence measure that quantifies the accuracy of hash generation and validation relative to an ideal, error-free scenario. The proposed confidence measure can serve as a principled heuristic for practical applications, independent of the underlying assumptions.
@@
$P(W_{\alpha})={1\over\sqrt{2\pi}\delta W_{\alpha}}\exp\left({-\frac{(W_{\alpha}-\bar{W}_{\alpha})^{2}}{2\delta W_{\alpha}^{2}}}\right).$ (11)
@@
$=$ ${1\over\sqrt{\pi}}\int_{d_{\alpha}}^{\infty}e^{-x^{2}}dx={1\over 2}[1-{\rm erf}(d_{\alpha})]$
@@
Note that $\delta W_{\alpha}$ and $\mathcal{N}_{\max}(\mathcal{N}_{\mathcal{H}})$ (a monotonically increasing function of the block length), should be determined one-time as a public rule of the protocol. Targeting some operational delay, efficiency and spoof-resistance tolerable to the community, this can be done as one time work from a model or empirical data (see for example Figure 5). Solutions based only on Hamming distance (not requiring these parameters) are also shown to be viable—although these are typically of lower efficiency and less resilient to spoofing. A compromise in rule complexity combined with one-time parameter tuning, would be essential in a best deployed solution.
@@ V.2 Redefining Chainwork
If the work remains constant, then one can assign $\pm 1$ to valid/invalid blocks, we call this basic +/- Chainwork. In the case of a tie, the first path evaluated is preferred, following the same rule as in Bitcoin. This adjustment effectively resolves the issue by turning a hard validation fork to a soft validation fork that is resolved subject to a delay. If a user falls behind by two blocks, the majority compute power of miners will produce a third block, causing the strongest chain to advance by three blocks in agreement with the majority.
@@
VI Demonstrated Implementation
Realization of a quantum blockchain requires access to a quantum computer capable of solving complex problems where quantum advantage has been demonstrated or is potentially achievable. The generated distribution must exhibit sufficient structure to allow observables with nontrivial expectation values that can be measured reproducibly across multiple QPUs. Therefore, techniques such as boson sampling or random circuit sampling which cannot be formulated as efficient decision problems, are not suitable for this application (see Appendix A). Coarse-grained boson sampling has been proposed as a means to implement blockchains [28, 29], but to our knowledge there is no demonstrated implementation of a blockchain using quantum computers (see Appendix A.1). Currently, Ref. [1] is the only experimental demonstration of beyond-classical computation that meets the requirement of efficient cross-validation. Therefore, we base our demonstration on the same protocol. Since Ref. [1] has already established agreement with quantum mechanics and the infeasibility of classical simulation, our primary focus here is to demonstrate that multiple QPUs can generate and validate hashes, ensuring the blockchain operates reliably with high efficiency and small transaction delays.
Figure 3: Operation of an example blockchain with $50$ miners using four generally-available Advantage and Advantage2 QPUs. The mining difficulty was set to $\mathcal{N}_{\rm zeros}=32$ with basic +/-1 Chainwork defined in (18). Orange blocks show soft (resolved) forks, these blocks are included in no strongest chains (have been rejected by all miners). Blue blocks are immutable, included in all strongest chains (agreed upon by all miners). Gray and black blocks are in contention: common to some (but not all) strongest chains. Only black blocks are candidates for maximum Chainwork (are being mined) and have the potential for further branching. Blocks are labelled by the order of broadcast with the genesis block on the left and most recently mined block on the right. Either of the paths terminating in black nodes might be extended, but the orange path is no longer a candidate for the strongest chain. Blocks 10, 11, 12 and 13 contain delayed transactions, that require further mining (fork resolution) in order to be confirmed. A majority of miners believe node 13 has maximum Chainwork, and in this demonstration all miners consolidate on this branch as the chain develops (see Figure 4).
@@ VI Demonstrated Implementation
Figure 5: Mining efficiency, i.e., the fraction of blocks joining the strongest chain. The Chainwork uses either the simple +/-1 weighting, or the confidence-based weighting with $\mathcal{N}_{\rm max}=1,$ 22 or 33 with a constant $\delta W_{\alpha}$ extracted from demonstration data. Confidence-weighted strongest chains are more efficient. The mining process is accelerated without impact on statistical outcomes as described in Appendix  D.3. Results are obtained with resampling from demonstration-parameterized witness distributions to accelerate analysis, each point an average on 16 chains of length 512-1024. Resampling statistics are in agreement with full-demonstration efficiencies reported in Figure 4 and appendices. Further details and statistics are contained in Appendix D.
Blockchains, whether our proposal or classical, operate according to rules agreed by a community. Ideally the rule set should be small and persistent, so that rule changes that might cause forking of the user base are infrequent. As with all practical blockchains, rules should be calibrated and agreed based on the capabilities of near-term participants, and revisited periodically by the community. We test several rule sets building upon the well-established Bitcoin rule set, but excluding some essential components: such as periodic updates of problem hardness, modelling of network desynchronization and participation of fraudulent actors. Relative to Bitcoin there are additional parameters; most importantly the choice of an ensemble of unitary evolutions. Our proof of concept demonstrates the feasibility of high efficiency and low delay blockchains in this context, but leaves open the possibility to further optimize protocol rules as a function of the deployed computing environment. We note that significant practical enhancements in performance are achievable with little overhead by employing the mitigations outlined in Appendix C.
The demonstrations are conducted on four D-Wave Advantage and Advantage2 QPUs. Each QPU comprises a network of coupled qubits that undergo continuous evolution governed by a time-dependent Hamiltonian
@@
The quantum problem we solve for hash generation is the coherent quenching of spin glasses through quantum phase transitions. We embed cubic lattices of size $4{\times}4{\times}4$ dimers (128 qubits) on each of four different QPUs in Canada and the United States; supplementary results for $18{\times}18$ dimerized Biclique models (72 qubits) are limited to Appendix D. The critical dynamics ensure that by adjusting a single parameter (annealing time) different QPUs can produce consistent results despite variations in their energy scales [1]. All qubit biases are set to zero, $h_{i}=0$, and couplers are selected as a function of the input message $J_{ij}=\Theta(\mathcal{M})$. This is done in two steps: First the message goes through a classical cryptographic hash, e.g., SHA-256 in our case. Then the classical hash is used as a seed in a pseudo-random number generator that generates the unitary-evolution parameters ($J$ only, in our proof of concept). The classical hash is also used as the cryptographic identifier of the block in the block header of the next block. In our proposal based on the Bitcoin block (Figure 1) the quantum hash is only used for PoQ. This way, we can choose ${\cal N_{H}}={\cal N}_{\rm zeros}$ because nonzero bits in the quantum hash have no significance (see Appendix C.4 for a mitigation of this constraint). Using a spin-glass Hamiltonian for quantum evolution ensures that classical collision resistance carries over to the quantum hash. It is extremely unlikely that two spin-glass Hamiltonians with entirely different randomly chosen parameters collide in the resulting high-dimensional distribution. Furthermore, the projection matrix $G(\mathcal{M})$ is also randomly generated based on the message $\mathcal{M}$, making collisions even more improbable.
@@
Fig. 3 presents the initial-stage operation of a blockchain using four generally-available D-Wave Advantage and Advantage2 QPUs, involving 50 miners with $\mathcal{N}_{\rm zeros}=32$. The visualization captures the dynamic evolution of the blockchain, highlighting block statuses and mining activity. Blocks on the (majority) strongest chain are arranged chronologically from left to right, with the genesis block on the far left and the most recently mined block on the far right. Different block colors indicate their status within the consensus process: Blue blocks (immutable) are included in all strongest chains and are universally agreed upon. Orange blocks (soft forks) have been rejected by all miners and are not included in any of the strongest chains. Gray and black blocks appear in some but not all strongest chains, reflecting temporary disagreement among miners. Only black blocks are actively being mined, meaning they have the potential for further branching. This illustrates how consensus emerges dynamically as blocks are proposed and either accepted or rejected by the network.
@@
In Fig. 4, the blockchain is shown after 219 block broadcasts, with the genesis block repositioned centrally for clarity. By this stage, approximately 70% of proposed blocks have become immutable (blue), aligning well with theoretical expectations based on the bootstrapping methodology (see Appendix D.3). The remaining blocks continue to be evaluated, with only those on the strongest chains persisting in the blockchain ledger. This simulation demonstrates the robustness of the blockchain framework, showcasing its ability to maintain stability and achieve miner consensus despite the probabilistic nature of quantum computation. We are able to use statistical properties of the Chainwork described in Section D.2 to eliminate failed mining events, and create the blockchain with only $50\times 219$ total QPU programmings. Blockchains with more miners, and over $1000$ block broadcasts, are shown in the appendices.